OpenStudy (anonymous):

If 3000 dollars is invested in a bank account at an interest rate of 9 % per year, compounded continuously. How many years will it take for your balance to reach 50000 dollars?

OpenStudy (anonymous):

50,000 = 3,000 e^.09t ->e^.09t = 50 / 3 => .09t = ln ( 50/3) => t = ln ( 50/3)/ .09 = 31.3 yrs